# Fulde-Ferrell-Larkin-Ovchinnikov States and Topological Bogoliubov Fermi Surfaces in Altermagnets: an Analytical Study

### v1

@@ I Introduction
Altermagnetism (AM) (Smejkal2022-1, ), a novel type of collinear magnetism beyond the traditional ferromagnetism (FM) and antiferromagnetism (AFM), has recently attracted considerable attention (Smejkal2022-2, ; Mazin2022, ; Bai2024, ; Song2025, ; Guo2025, ; Jungwirth2025-1, ; Jungwirth2025-2, ). As a hybrid of FM and AFM, AM features compensated collinear spin configurations in real space (akin to conventional AFM) while exhibits non-relativistic spin splitting (NRSS) in momentum space (reminiscent of FM) (Lee2024, ; Souma2024, ; Reimers2024, ; Ding2024, ; Krempasky2024, ; Yang2025, ; Leiviska2024, ; Han2024, ; Zhou2025, ; Jiang2025, ; Zhang2025, ). Combing both merit of FM and AFM, AM offers exciting opportunities for novel applications in spintronics (Naka2019, ; Smejkal2020, ; Reichlova2024, ), quantum geometry (Fang2024, ; Gong2024, ; LiY2024-1, ; Hernandez2025, ; Liu2025, ), and superconductivity (Mazin2022-2, ; Zhu2023, ; Brekke2023, ; Bose2024, ; Wu2025, ; LiY2024-2, ; deCarvalho2024, ; Chakraborty2024-2, ; LiY2023, ; Ghorashi2024, ; Hodge2025, ; Ouassou2023, ; Lu2024, ; Fukaya2025, ; Banerjee2024, ; Chakraborty2024-3, ; Papaj2023, ; Sun2023, ; Das2024, ; Wei2024, ; Carvalho2024, ; Parthenios2025, ).
@@
AM also provides a promising platform for the FFLO state for two main reasons. First, FFLO phases typically require a nonzero spin imbalance, which is usually achieved via external Zeeman fields. However, in conventional superconductors, the orbital effect induced by magnetic fields can suppress the superconductivity before the FFLO emerges (Maki1966, ; Gruenberg1966, ). In contrast, AM naturally hosts NRSS without the need of external magnetic field, thereby avoiding detrimental orbital effects. Second, conventional FFLO state involves spontaneously breaking of the continuous $U(1)$ rotational symmetry of the Fermi surfaces, which, according to Nambu-Goldstone theorem (Nambu1960, ; Goldstone1961, ; Goldstone1962, ), leads to gapless modes associated with fluctuations between equivalent $\bf q$}. In AM, the continuous rotational symmetry is reduced to a discrete one Wu2007 , significantly suppressing such fluctuations. These unique advantages have motivated several theoretical investigations of FFLO phases in altermagnetic metals (Feiguin2009, ; Soto-Garrido2014, ; Gukelberger2014, ; Sumita2023, ; Chakraborty2024-1, ; ZhangS2024, ; Hong2025, ; Hu2025-1, ; Hu2025-2, ).
@@
Despite these efforts, the existence of FFLO state in two-dimensional (2D) $d$-wave AM Fermi gas with $s$-wave pairing remains debated. Using a continuum model, Soto-Garrido and co-workers (Soto-Garrido2014, ) reported the absence of FFLO state in a 2D $d_{x^{2}-y^{2}}$-wave AM Fermi gas. This conclusion is corroborated by Feiguin et al. (Feiguin2009, ) and Chakraborty et al. (Chakraborty2024-1, ) through a square lattice model. Nevertheless, more recent studies challenge this result. Investigating a similar lattice model to that of Chakraborty et al. (Chakraborty2024-1, ), Hong and co-workers (Hong2025, ) found that FFLO state does emerge under purely $s$-wave pairing. Subsequently, Hu and co-workers (Hu2025-1, ) conducted a numerical calculation on the continuum model close to Soto-Garrido et al. (Soto-Garrido2014, ) and confirmed the existence of FFLO state, aligning with the findings of Hong et al. (Hong2025, ). These conflicting reports raise a fundamental question: can 2D $d$-wave AM Fermi gas supports the FFLO state driven solely by $s$-wave pairing?
@@
In this work, we give a definitive and affirmative answer by analytically determining the ground-state phase diagram. We consider two physically relevant situations: (i) fixed $\mu/\varepsilon_{B}$, where $\mu$ is the total chemical potential and $\varepsilon_{B}$ is the two-body binding energy which is experimentally controllable in the ultracold atom systems; and (ii) fixed total particle number $N$, which is appropriate for describing electronic superconductors. In both situations, we demonstrate that FFLO phases do indeed exist. What’s more, another exotic unconventional form of superconductivity - nodal SF hosting topological Bogoliubov Fermi surface (TBFS) - also emerges in both situations.
@@
The rest of the paper is organized as follows. In the next section (Sec. II), we outline the model Hamiltonian that describes a $d$-wave altermagnetic metal. In Sec. III, we briefly discuss how to solve the model Hamiltonian by using the standard Bogoliubov de-Gennes (BdG) mean-field theory, with an inhomogeneous FF order parameter. In Sec. IV, we present the phase diagram between normal (N) phase, nodal SF, and SF. The nontrivial $\mathbb{Z}_{2}$ topology of BFS is characterized by the Pfaffian of the BdG Hamiltonian. In Sec. V, we discuss in details the FFLO phase boundary driven by the NRSS. Finally, we conclude in Sec. VI and present some outlooks for future works.
@@ II Model Hamiltonian
For the non-interacting part ${\cal H}_{0}$, $\hat{\varepsilon}_{{\bf k}}\equiv-{\bf\nabla}^{2}$ is the operator for kinetic energy where $\hat{k}_{x}=-i\partial_{x}$, $\hat{k}_{y}=-i\partial_{y}$. $\hat{\xi}_{{\bf k}}\equiv\hat{\varepsilon}_{{\bf k}}-\mu$ is the modified kinetic energy operator where $\mu$ is the chemical potential. The chemical potential is tuned to yield the given total particle density $n=n_{\uparrow}+n_{\downarrow}$. $\hat{J}_{\mathbf{k}}\equiv t_{\textrm{AM}}2\hat{k}_{x}\hat{k}_{y}$ where $s(\uparrow)=+1$ and $s(\downarrow)=-1$ is the $d_{xy}$-wave altermagnetic spin splitting, whose strength is controlled by the dimensionless parameter $t_{\textrm{AM}}$ (without the loss of generality, $t_{\textrm{AM}}$ is set to be positive here). In 2D, there are two independent $d$-wave spin splitting: $d_{xy}$ and $d_{x^{2}-y^{2}}$. As $d_{x^{2}-y^{2}}$ can be obtained from $d_{xy}$ by $\pi/4$ rotation, the results obtained here can be generalized to $d_{xy}$-wave altermagnetic spin splitting straightforwardly. Without $\hat{J}_{\mathbf{k}}$, the spin degenerated Fermi surface is a circle with radius $\sqrt{\mu}$, which conserves $U(1)$ rotational symmetry (see the dashed circle in Fig. 1(a)). The introduction of $\hat{J}_{\mathbf{k}}$ distorts the Fermi surface into two orthogonal ellipses (denoted as Fermi ellipse here) with equal semi-major axis $a\equiv\sqrt{\mu/(1-t_{\textrm{AM}})}$ and semi-minor axis $b\equiv\sqrt{\mu/(1+t_{\textrm{AM}})}$ (see Fig. 1(a)), in this regard, $t_{\textrm{AM}}$ also measures the eccentricity of the Fermi ellipses at a given $\mu$. These two Fermi ellipses together exhibit $d_{xy}$-wave form (see Fig. 1(a)): spin up and down ellipse has major axis along $\left(1,-1\right)$ and $\left(1,1\right)$ direction, respectively. Since the two Fermi ellipses always have the same area even with finite momentum boost in the FFLO phase, the system remains spin-population balanced in the whole phase diagram. Such a circle-to-ellipse distortion breaks the continuous $U(1)$ to the discrete $C_{4z}$ symmetry, leaving four degeneracies along both $(1,0)$ and $(0,1)$ directions. As mentioned before, such continuous symmetry breaking would stabilize the FFLO state.
@@
$\frac{1}{g}=-\frac{1}{\mathcal{A}}\sum_{\mathbf{k}}\frac{1}{2\varepsilon_{{\bf k}}+\varepsilon_{B}},$ (4)
@@ III Mean-field theory
To search for the FF superconductivity, we assume a FF-like order parameter $\Delta({\bf x})=-g\left\langle\psi_{\downarrow}\left({\bf x}\right)\psi_{\uparrow}\left({\bf x}\right)\right\rangle=\Delta\exp[i\mathbf{q}\cdot\mathbf{x}]$ at positive $\mathbf{x}$ with the center-of-mass momentum $\mathbf{q}$, so $\left|\Delta({\bf x})\right|^{2}=\Delta^{2}$ (below we take $\Delta$ positive) and consider the standard mean-field decoupling of the interaction Hamiltonian density (Hu2006, ; Liu2013, ),
@@
Figure 1: (a) $d_{xy}$-wave spin-splitting by plotting $\xi_{{\bf k}}+s(\sigma)J_{{\bf k}}=0$, here $\mu=1$. $t_{\textrm{AM}}=0$ gives the dashed circle, and $t_{\textrm{AM}}=0.5$ gives the red and blue ellipses, corresponding to spin up and down channel, respectively. (b) Contour plot of $E_{{\bf k\pm}}=0$, here $\mu=1$, $t_{\textrm{AM}}=0.5$, and $\Delta=0$. The dash line corresponds to $\sum_{\eta}E_{{\bf k\eta}}=0$. Cyan and blue colored regions correspond to negative $E_{{\bf k+}}$ and $E_{{\bf k-}}$. (c) Curve for $E_{k}$ at different $t_{\textrm{AM}}$ values, here $\mu=1,\Delta=0.2$. (d) Plot of $\bar{t}_{\textrm{AM}}$ (see Eq. (42) for definition) as a function of $\Delta/\mu$. (e)-(f) Anisotropic Bogoliubov quasiparticle energy $E_{+}(\mathbf{k})$ and $E_{-}(\mathbf{k})$. The while solid line represent the Bogoliubov Fermi surface. Here $\mu=1$, $\Delta=0.2$, and $t_{\textrm{AM}}=0.5$, which corresponds to the red dot in (d).
@@ III.1 Bogoliubov quasiparticles
$\left[{\cal H}_{\textrm{BdG}}\right]$ can be diagonalized with eigenvectors $u_{{\bf k}\uparrow}^{(\eta)},v_{{\bf k}\downarrow}^{(\eta)}]^{T}$ and corresponding eigenvalues $\eta E_{{\bf k}\eta}$, labeled by the branch index $\eta=\pm$. Explicitly, we have the Bogoliubov quasiparticle energy spectrum,
@@
III.2 Thermodynamic potential and ground state energy

### v2

@@ I Introduction
Altermagnetism (AM) (Smejkal2022-1, ), a novel type of collinear magnetism beyond the traditional ferromagnetism (FM) and antiferromagnetism (AFM), has recently attracted considerable attention (Smejkal2022-2, ; Mazin2022, ; Bai2024, ; Song2025, ; Guo2025, ; Jungwirth2025-1, ; Jungwirth2025-2, ). As a hybrid of FM and AFM, AM features compensated collinear spin configurations in real space (akin to conventional AFM) while exhibits non-relativistic spin splitting (NRSS) in momentum space (reminiscent of FM) (Lee2024, ; Souma2024, ; Reimers2024, ; Ding2024, ; Krempasky2024, ; Yang2025, ; Leiviska2024, ; Han2024, ; Zhou2025, ; Jiang2025, ; Zhang2025, ). Combing both merit of FM and AFM, AM offers exciting opportunities for novel applications in spintronics (Naka2019, ; Smejkal2020, ; Reichlova2024, ), quantum geometry (Fang2024, ; Gong2024, ; LiY2024-1, ; Hernandez2025, ; Liu2025, ; Antonenko2025, ), and superconductivity (Mazin2022-2, ; Zhu2023, ; Brekke2023, ; Bose2024, ; Wu2025, ; Maeda2025, ; LiY2024-2, ; Chakraborty2024-2, ; LiY2023, ; Ghorashi2024, ; Mondal2025, ; Hodge2025, ; Ouassou2023, ; Lu2024, ; Fukaya2025, ; Zhao2025, ; Banerjee2024, ; Chakraborty2024-3, ; Papaj2023, ; Sun2023, ; Das2024, ; Wei2024, ; Carvalho2024, ; Parthenios2025, ; Fukaya2025-JPCM, ; Liu2025-2, ).
@@
AM also provides a promising platform for the FFLO state for two main reasons. First, FFLO phases typically require a nonzero spin imbalance, which is usually achieved via external Zeeman fields. However, in conventional superconductors, the orbital effect induced by magnetic fields can suppress the superconductivity before the FFLO emerges (Maki1966, ; Gruenberg1966, ). In contrast, AM naturally hosts NRSS without the need of external magnetic field, thereby avoiding detrimental orbital effects. Second, conventional FFLO state involves spontaneously breaking of the continuous $U(1)$ rotational symmetry of the Fermi surfaces, which, according to Nambu-Goldstone theorem (Nambu1960, ; Goldstone1961, ; Goldstone1962, ), leads to gapless modes associated with fluctuations between equivalent $\bf q$}. In AM, the continuous rotational symmetry is reduced to a discrete one Wu2007 , significantly suppressing such fluctuations. These unique advantages have motivated several theoretical investigations of FFLO phases in altermagnetic metals (Feiguin2009, ; Soto-Garrido2014, ; Gukelberger2014, ; Sumita2023, ; Chakraborty2024-1, ; ZhangS2024, ; Hong2025, ; Hu2025-1, ; Hu2025-2, ; Sim2025, ; Sumita2025, ).
@@
Despite these efforts, the existence of a FFLO state in a two-dimensional (2D) $d$-wave AM Fermi gas with purely $s$-wave pairing remains controversial. Numerical simulations on a 2D square lattice by Feiguin et al. (Feiguin2009, ), Gukelberger et al. (Gukelberger2014, ), and Chakraborty et al. (Chakraborty2024-1, ) reported no evidence of a FFLO state within the parameter regimes explored. This absence was further supported by the analytical analysis of the corresponding continuum model according to Soto-Garrido et al. (Soto-Garrido2014, ). However, more recent studies challenge this conclusion. Using a lattice model closely related to that of Chakraborty et al. (Chakraborty2024-1, ), Hong et al. (Hong2025, ) found that a FFLO state does emerge under purely $s$-wave pairing. Consistent with this, Hu et al. (Hu2025-1, ) conducted a numerical calculation of the continuum model close to that of Soto-Garrido et al. (Soto-Garrido2014, ) and confirmed the presence of a FFLO state. These conflicting results raise a fundamental question: Can a 2D $d$-wave AM Fermi gas support a FFLO state driven solely by $s$-wave pairing?
Figure 1: (a) Zero temperature phase diagram at fixed chemical potential $\mu$ (normalized to the two-body binding energy $\varepsilon_{B}$) versus non-relativistic spin splitting $t_{\textrm{AM}}$. The solid black/dashed black/solid red curve giving the phase boundary is determined by Eq. (37)/Eq. (42)/Eq. (52), respectively. (b) Zero temperature phase diagram at fixed total particle number NN as a function of the binding energy $\varepsilon_{B}$ normalized to the Fermi energy $\varepsilon_{F}$. The solid black/solid red curve is given by Eq. (38)/Eq. (53), respectively. The black curves in both figures show the Pauli-limit, upper altermagnetic coupling strength, where a SF with zero center-of-mass momentum $\mathbf{q}=0$ becomes unstable towards a normal state, when the possibility of an inhomogeneous FFLO is not considered. In the weak-coupling regime (i.e., for large $\mu/\varepsilon_{B}$ or small $\varepsilon_{B}/\varepsilon_{F}$), the solid black curve in (a) and (b) provides an excellent estimate for the phase boundary separating a TBFS phase and a FFLO state at fixed $\mu$, and a BCS SF phase and a FFLO state at fixed $N$, respectively.
I.1 Summary of main results
@@
In this work, we give a definitive and affirmative answer by analytically determining the ground-state phase diagram at $T$ = 0 K. We consider two physically relevant settings: (i) fixed $\mu/\varepsilon_{B}$, where $\mu$ is the total chemical potential and $\varepsilon_{B}$ is the two-body binding energy which is experimentally controllable in the ultracold atom systems; and (ii) fixed total particle number $N$, which is appropriate for describing electronic superconductors.
Fig.1(a) shows the phase diagram in the fixed-chemical-potential ensemble. We find a finite window of FFLO stability that expands with increasing $d$-wave altermagnetic spin splitting $t_{AM}$ (a dimensionless parameter defined below). Adjacent to the FFLO state is a nodal SF hosting topological Bogoliubov Fermi surfaces (TBFSs), whose region of stability is comparatively narrow.
Fig.1(b) shows the phase diagram at fixed total particle number. Here, the horizontal axes is normalized to the Fermi energy $\varepsilon_{F}$ (defined below), which is determined by the imposed particle number. In this ensemble, we again find that the FFLO window broadens with increasing $t_{AM}$, appearing between the conventional BCS SF and the normal altermagnetic metallic phase.
To obtain the FFLO-normal phase boundary (red curves in Fig.1), one must evaluate the two polar integration integrals $I_{3}$ and $I_{4}$ defined in Eq. (48). The central technical contribution of this work is an analytical evaluation of $I_{3}$ and $I_{4}$, which, to our best knowledge, has not been carried out exactly in previous works (Soto-Garrido2014, ; ZhangS2024, ). In general, two approaches exist for polar integration in 2D. The first integrates over the radial distance $k$ followed by the polar angle $\theta$. This integration order is widely employed in evaluating 2D polar integration, as $I_{1}$ and $I_{2}$ in this work. The second integrates over the polar angle $\theta$ first, which is recast to a complex integral on the unit circle $|z|=1$ in the complex $z$ plane, and subsequently performs the remaining $k$-integration. This second procedure is used to evaluate $I_{3}$ and $I_{4}$. After performing the $\theta$-integration, the resulting expressions depend on both $k$ and the FFLO center-of-mass momentum $\bf{q}$, and acquire different analytic forms in distinct ($k$, $\bf{q}$) domains. The boundaries between these domains determine the critical momentum $\bf{q}_{c}$ marking the onset of the FFLO state. For $\bf{q}$ oriented along the (1, 1) direction, we found that $|\bf{q}_{c}|$ admits a simple geometric interpretation: it is the momentum at which the spin-up and spin-down Fermi surfaces become nested (see Fig.4(a)).
@@
The remainder of the paper is organized as follows. In the next section (Sec. II), we outline the model Hamiltonian that describes a $d$-wave altermagnetic metal. In Sec. III, we briefly discuss how to solve the model Hamiltonian by using the standard Bogoliubov de-Gennes (BdG) mean-field theory, with an inhomogeneous FF order parameter. In Sec. IV, we present the phase diagram between normal, nodal SF, and BCS SF phases. The nontrivial $\mathbb{Z}_{2}$ topology of BFS is characterized by the Pfaffian of the BdG Hamiltonian. In Sec. V, we discuss in details the FFLO phase boundary driven by the NRSS. Finally, we conclude in Sec. VI and present some outlooks for future works.
@@ II Model Hamiltonian
For the non-interacting part ${\cal H}_{0}$, $\hat{\varepsilon}_{{\bf k}}\equiv-{\bf\nabla}^{2}$ is the operator for kinetic energy where $\hat{k}_{x}=-i\partial_{x}$, $\hat{k}_{y}=-i\partial_{y}$. Operating on a plane wave with wave vector $\bf{k}$, it gives the kinetic energy $\varepsilon_{\bf{k}}=k^{2}$, where $k$ is the length of $\bf{k}$. $\hat{\xi}_{{\bf k}}\equiv\hat{\varepsilon}_{{\bf k}}-\mu$ is the modified kinetic energy operator where $\mu$ is the chemical potential. The chemical potential is tuned to yield the given total particle density $n=n_{\uparrow}+n_{\downarrow}$. $\hat{J}_{\mathbf{k}}\equiv t_{\textrm{AM}}2\hat{k}_{x}\hat{k}_{y}$ where $s(\uparrow)=+1$ and $s(\downarrow)=-1$ is the $d_{xy}$-wave altermagnetic spin splitting, whose strength is controlled by the dimensionless parameter $t_{\textrm{AM}}$ (without the loss of generality, $t_{\textrm{AM}}$ is set to be positive here). In 2D, there are two independent $d$-wave spin splitting: $d_{xy}$ and $d_{x^{2}-y^{2}}$. As $d_{x^{2}-y^{2}}$ can be obtained from $d_{xy}$ by $\pi/4$ rotation, the results obtained here can be generalized to $d_{xy}$-wave altermagnetic spin splitting straightforwardly. Without $\hat{J}_{\mathbf{k}}$, the spin degenerated Fermi surface is a circle with radius $\sqrt{\mu}$, which conserves $U(1)$ rotational symmetry (see the dashed circle in Fig. 2(a)). The introduction of $\hat{J}_{\mathbf{k}}$ distorts the Fermi surface into two orthogonal ellipses (denoted as Fermi ellipse here) with equal semi-major axis $a\equiv\sqrt{\mu/(1-t_{\textrm{AM}})}$ and semi-minor axis $b\equiv\sqrt{\mu/(1+t_{\textrm{AM}})}$ (see Fig. 2(a)), in this regard, $t_{\textrm{AM}}$ also measures the eccentricity of the Fermi ellipses at a given $\mu$. These two Fermi ellipses together exhibit $d_{xy}$-wave form (see Fig. 2(a)): spin up and down ellipse has major axis along $\left(1,-1\right)$ and $\left(1,1\right)$ direction, respectively. Since the two Fermi ellipses always have the same area even with finite momentum boost in the FFLO phase, the system remains spin-population balanced in the whole phase diagram. Such a circle-to-ellipse distortion breaks the continuous $U(1)$ to the discrete $C_{4z}$ symmetry, leaving four degeneracies along both $(1,0)$ and $(0,1)$ directions. As mentioned before, such continuous symmetry breaking would stabilize the FFLO state.
@@
$\frac{1}{g}=-\frac{1}{\mathcal{A}}\sum_{\mathbf{k}}\frac{1}{2\varepsilon_{\mathbf{k}}+\varepsilon_{B}},$ (4)
@@
Figure 2: (a) Schematic $d_{xy}$-wave spin splitting obtained by plotting $\xi_{\bf{k}}+s(\sigma)J_{\bf{k}}=0$. Here $\mu=1$. $t_{\textrm{AM}}=0$ gives the dashed circle, and $t_{\textrm{AM}}=0.5$ gives the red and blue ellipses, corresponding to spin-up and spin-down channels, respectively. The semi-major and semi-minor of each ellipse is denoted as $a$ and $b$. (b) Contour plot of $E_{\bf{k}\pm}=0$ (see Eq. (12) for definition) at normal phase, here $\mu=1$, $t_{\textrm{AM}}=0.5$, and $\Delta=0$. The dash line corresponds to $\sum_{\eta}E_{\bf{k}\eta}=0$. Cyan and blue colored regions correspond to negative $E_{\bf{k}+}$ and $E_{\bf{k}-}$. (c) Plot of $E_{\bf{k}}$ (see Eq. (40) for definition) at different $t_{\textrm{AM}}$ values, here $\mu=1,\Delta=0.2$. When $t_{AM}$ is small, $E_{\bf{k}\pm}$ is always positive, indicating a BCS SF phase. When $t_{AM}$ is large, $E_{\bf{k}\pm}$ will vanishing, meaning the emergence of Bogoliubov Fermi surfaces. (d) Plot of $\bar{t}_{\textrm{AM}}$ (see Eq. (41) for definition) as a function of $\Delta/\mu$. (e)-(f) Anisotropic Bogoliubov quasiparticle energy $E_{\bf{k}+}$ and $E_{\bf{k}-}$ at BCS SF phase. The white solid lines represent the Bogoliubov Fermi surfaces $E_{\bf{k}\pm}=0$, which separate the whole $\bf{k}$ domain into two regions labelled by $\bf{k}_{\pm}$. Here $\mu=1$, $\Delta=0.2$, and $t_{\textrm{AM}}=0.5$, which corresponds to the red dot in (d).
@@ III Mean-field theory
To search for the FFLO state, for simplicity, we employ a plane-wave ansatz for the pairing order parameter $\Delta({\bf x})=-g\left\langle\psi_{\downarrow}\left({\bf x}\right)\psi_{\uparrow}\left({\bf x}\right)\right\rangle=\Delta e^{i\mathbf{q}\cdot\mathbf{x}}$ at position $\mathbf{x}$ with the center-of-mass momentum $\mathbf{q}$. This is equivalent to the phase-modulated FF form introduced by Fulde and Ferrel (Fulde1964, ), here we still refer to the resulting state as the FFLO state. The amplitude-modulated LO type, will be explored in future work.
Under plane-wave ansatz, $\left|\Delta({\bf x})\right|^{2}=\Delta^{2}$ (below we take $\Delta$ positive), and consider the standard mean-field decoupling of the interaction Hamiltonian density (Hu2006, ; Liu2013, ),
@@ III.1 Bogoliubov quasiparticles
$\left[{\cal H}_{\textrm{BdG}}\right]$ can be diagonalized with eigenvectors $u_{{\bf k}\uparrow}^{(\eta)},v_{{\bf k}\downarrow}^{(\eta)}]^{T}$ and corresponding eigenvalues $\eta E_{{\bf k}\eta}$, labelled by the branch index $\eta=\pm$. Explicitly, we have the Bogoliubov quasiparticle energy spectrum,
@@
III.2 Thermodynamic potential and ground-state energy